OpenStudy (anonymous):

Evaluate without using a calculator arcsin1/2

OpenStudy (anonymous):

sin(x) = 1/2 think of a number when plugged in to the sin function would give you 1/2?

OpenStudy (anonymous):

one of those trig values that you need to memorize. arcsin(1/2) = pi/6